In R Studio, Does the number of tadpoles vary between the relative altitudes at which the sites were located? If so, which of the altitude level means are significantly different?

Site Rain Temp Altitude Tadpoles Breeding
1 769.77 22.37 Low 97 61
2 664.85 19.46 Low 86 70
3 807.13 23.17 Low 104 61
4 620.05 17.85 Low 117 113
5 673.21 20.29 Low 98 92
Answer #1

solution:

   Here the statistical hypothesis to be tested is as follows,

Null hypothesis: μ = 48

Alternative hypothesis: μ ≠ 48

Where, μ = mean number of breeding ponds ,

R - code is as follows;

Mu= 48 # True or given mean
Breeding=c(61,70,61,113,92,100,102,66,64,18)
Xbar= mean(Breeding) # sample mean
Xbar
SE=sd(Breeding) # Sample standard deviation/ standard error
SE

# Here test statistic computed as follows,
  
Tcal= (Xbar- Mu)/SE
Tcal
Ttab=qt(0.975,9) # Tabulated value of t-probability distribution at 9 d.f. and 95% level of significance.
Ttab

# Since Tcal < Ttab , we may accept null hypothesis.


R-code out put ;-

> Mu= 48       # True or given mean
> Breeding=c(61,70,61,113,92,100,102,66,64,18)
> Xbar= mean(Breeding)      # sample mean 
> Xbar
[1] 74.7
> SE=sd(Breeding)           # Sample standard deviation/ standard error 
> SE
[1] 27.79708
> 
> # Here test statistic computed as follows,
>   
> Tcal= (Xbar- Mu)/SE
> Tcal
[1] 0.9605325
> Ttab=qt(0.975,9)     # Tabulated value of t-probability distribution at 9 d.f. and 95% level of significance.
> Ttab 
[1] 2.262157
> 
> # Since Tcal < Ttab , we may accept null hypothesis.

Hence we may claim that the mean number of breeding ponds found in NSW is same as the mean number found in the Victorian surveys.

please give me thumb up